solve for x: radical sign x-4= x-10 (the sign is over x-4 only)I got x=13 x=-8, but the -8 was extraneous....am i correct???
Found 2 solutions by stanbon, asha:
Answer by stanbon(75887)

radical sign x-4= x-10 (the sign is over x-4 only)I got x=13 x=-8, but the -8 was extraneous....am i correct???
-----------
sqrt(x-4) = x-10
Square both sides to get:
x-4=x^2-20x+100
x^2-21x+104=0
(x-13)(x-8)=0
x=13 or x=8
Cheers,
Stan H.

Answer by asha(30)

radical sign x-4 =x-10
squaring both the sides we get
x-4 = (x-10)^2
x-4 = x^2-20x+100
taking all the terms to one side,we get
x^2-20x-x+100+4
x^2-21x+104
splitting the middle term so that sum is -21 and product is 104
we get the factors -8 and -13
x^2-8x-13x+104=0
x(x-8)-13(x-8)=0
(X-8)(x-13)=0
x=8 and x=13 are the solutions
good luck!!!